a pail and a jug contain q litres of water. The pail contains 3 times as much water as the jug. Find the volume of water in the jug in terms of q?
Answer by Theo(13342)

let p represent the number of liters of water in the pail.
let j represent the number of liters of water in the jug.
p + j = q
the pail contains 3 times as much water as the jog.
p = 3j
replace p with 3j in the formula of p + j = q to get:
3j + j = q
combine like terms to get:
4j = q
divide both sides of the equation by 4 to get:
j = q/4

that's your solution.
the volume of water in the jug is equal to the total volume of water / 4.

let's assume that j = 2 liters.
then p = 3*j = 6 liters.
q would then be equal to 8 liters.
you would get:
p + j = 8
p = 3j
formula becomes:
3j + j = 8 which becomes:
4j = 8 which becomes:
j = 2
since p = 3j, then p = 6
since 2 = 8/4, and j = 2 while q = 8, this means that:
j = q/4

formula checks out with number so the solution looks good.
the solution is:
j = q/4.